OpenStudy (shesolitt):

I WILL MEDAL AND FAN Derive the equation of the parabola with a focus at (2, −1) and a directrix of y = −one half. f(x) = −(x + 2)2 − three fourths f(x) = (x + 2)2 + three fourths f(x) = −(x − 2)2 + three fourths f(x) = −(x − 2)2 − three fourths

OpenStudy (shesolitt):

@Afro @563blackghost @CallMeKiki

OpenStudy (shesolitt):

@mathstudent55 @ShadowLegendX

OpenStudy (daniyil_the_spy):

ok

OpenStudy (daniyil_the_spy):

sry I don't know

OpenStudy (dumbcow):

 From this we know that p = 1/4 (distance from vertex to directrix) We know that parabola is negative and opens down vertex = (2,-3/4) Equation: \[y =- \frac{1}{4p}(x-h) + k\] plugging in values you get: \[y = -(x-2) - \frac{3}{4}\]